Scores on the common final exam given in a large enrollment course were normally distributed with mean of 69.35 and standard deviation of 12.93. The department has the rule that in order to receive an A in the course a student’s score must be in the top 10% of all exam scres. Find the minium exam score that meets this requiremnt..
Answer #1

Solution :

Given that,  

mean = = 69.35

standard deviation = = 12.93

Using standard normal table ,

P(Z > z) = 10%

1 - P(Z < z) = 0.1

P(Z < z) = 1 - 0.1

P(Z < 1.28) = 0.9

z = 1.28

Using z-score formula,

x = z * +

x = 1.28 * 12.93 + 69.35 = 85.90

Minimum exam score = 85.90